# Score Change of Variables

### v1

@@ 5 Discussion
This work introduces a change of variables formula for score functions, providing a principled approach to adapt score-based methods across different spaces. By enabling transformations that accommodate specific data geometries and constraints, our result increases flexibility in model design and training. We demonstrate this through a reverse-time Itô lemma for transformed diffusion models and a generalized sliced score matching framework. Together, these contributions lay a mathematical foundation for transforming score-based algorithms, opening up new avenues for research and applications.
@@ 5.1 Related Work
Our approach complements existing research on distribution transformations, such as normalizing flows [20, 8], which focus on mapping densities rather than their scores. It also connects to recent developments in score-based diffusion models on constrained domains and manifolds [6, 4, 12, 18, 11]. In particular, [4] uses a change of variables for biological sequence generation, and our framework provides a unifying perspective and rigorous analytical tools to understand such transformations more broadly.
@@ 5.2 Limitations
A primary limitation is the computational cost of computing Jacobians and potential numerical instabilities when applying complex transformations. Our theoretical analysis ensures the existence of transformed score functions but does not guarantee numerical stability or efficient convergence for sampling. Moreover, the requirement of smooth, well-conditioned transformations restricts the class of applicable mappings. Finally, current variance reduction techniques for GSSM are limited to specific families of transformations.
@@ 5.3 Future Directions
Several promising directions can further advance this line of work.
First, while we focused on analytically tractable mappings, exploring more expressive transformations, such as invertible neural networks, could offer greater flexibility and adaptability to complex data geometries. This data-driven approach could lead to significant performance improvements, but also increase computational costs.
Second, deeper theoretical study is needed to understand the interplay between transformations and diffusion processes. Such insights could guide the selection of optimal transformations, challenging the prevailing assumption that the same space should be used for both training and sampling. This may lead to improved sampling efficiency and better-tailored geometries for specific applications.
Third, extending the score change of variables framework to other score-based methods, including variational inference [19] and energy-based models [24], presents a compelling direction for future research and could unlock new capabilities within these frameworks.
In summary, the theoretical foundations established here open the door to more versatile and effective score-based methodologies, offering opportunities for improved performance, stability, and interpretability across a range of applications.

### v2

@@ 5 Discussion
This work explores a change of variables formula for score functions, providing a principled approach to adapt score-based methods across different spaces. By enabling transformations that accommodate specific data geometries and constraints, our result increases flexibility in model design and training. We demonstrate this through a reverse-time Itô lemma for transformed diffusion models and a generalized sliced score matching framework. Together, these contributions lay a mathematical foundation for transforming score-based algorithms, opening up new avenues for research and applications.
@@ 5.1 Related Work
Our work on score function transformations intersects with several important areas of research in machine learning and statistics. In the realm of distribution transformations, our approach complements existing research such as normalizing flows [25, 9], which primarily focus on mapping densities rather than their scores. A review by [21] highlights the importance and breadth of change of variable formulas in generative modeling. While these works deal mainly with density transformations, our focus on score function transformations offers a perspective particularly useful in generative diffusion scenarios.
Recent advancements in score-based diffusion models have challenged the conventional use of Gaussian forward and reverse processes in generative modeling. Several studies have explored the application of diffusion models to constrained domains where standard Gaussian diffusion is inadequate. For instance, [7] adapted diffusion processes to manifold structures. Other researchers have investigated the use of reflected Brownian motion for sampling in constrained domains [22, 13], further expanding the applicability of diffusion models to complex geometries.
Our chess example, which explores diffusion on the probability simplex by transforming a VP SDE, aligns closely with the work of [14]. However, while they derive the score and reverse-time process entirely within the more complex simplex space, our approach offers a simpler model that is easier to train and sample from. In a similar vein to our paper, [4] employs a change of variables to train a score model with a Jacobi forward diffusion process for biological sequence generation. Our reverse-time Itô lemma generalizes and expands upon these efforts, providing a more comprehensive framework for score transformations.
The field has also witnessed the development of more exotic diffusion processes aimed at improving sample quality. Notable examples include the generative fractional diffusion models introduced by [23], which extend the concept of diffusion to fractional dynamics, and the critically-damped Langevin diffusion proposed by [10] for score-based generative modeling, offering enhanced sampling efficiency. The score change of variables framework can potentially be applied to sampling on these diverse domains and processes, offering significant benefits in cases like the simplex diffusion of [14]. While for some scenarios, no straightforward transformation may exist, we believe this avenue of research—finding transformations to simplify training and sampling in these complex domains—remains largely unexplored and holds considerable promise.
Score matching techniques have been instrumental in estimating and learning score functions, with various approaches developed over the years. Our work on generalized sliced score matching (GSSM) builds upon the original sliced score matching [29], extending the concept to allow for more flexible, non-linear projections in the score estimation process. This approach relates to other established methods in the field, such as the foundational score matching [18], denoising score matching [32], and the Stein spectral gradient estimator [28]. Each of these techniques offers unique advantages in different contexts, and the score change of variables framework presented in this paper may open avenues for developing new score matching methods. For instance, it might enable the adaptation of existing techniques to transformed spaces, potentially simplifying score estimation in complex domains or for intricate probability distributions.
By examining score function transformations and their applications in diffusion models and score matching, this work aims to contribute to the ongoing dialogue in these fields. The framework presented here offers potential enhancements to a range of machine learning tasks involving score functions and generative modeling, bridging theoretical concepts with practical applications. As the field continues to evolve, further investigation into the interplay between these areas may yield additional insights and methodologies for addressing complex problems in score-based machine learning.
@@ 5.2 Limitations
While our framework provides a general approach for transforming score functions, several limitations warrant consideration:
  • Computational and Numerical Challenges: The computational cost of Jacobian and Hessian calculations, particularly in the reverse-time Itô lemma, can be significant for complex transformations where analytical solutions are not tractable. While we prove the existence of transformed score functions, our framework does not guarantee numerical stability or efficient convergence for sampling in transformed spaces.
  • Restrictive Transformation Requirements: Our analysis is currently limited to smooth, invertible mappings, excluding a wide class of potentially useful non-differentiable or non-invertible transformations. Exploring the reverse-time Itô lemma with non-invertible transformations, akin to the original Itô lemma, represents an interesting direction for future work.
  • Limited Variance Reduction Techniques: For Generalized Sliced Score Matching (GSSM), current variance reduction techniques are specific to certain families of transformations and often require individual derivation. This limits their general applicability and increases the computational burden for novel transformations.
  • Practical Utility: While we demonstrate potential applications, such as controlling sampling dynamics in the chess example, the broader practical benefits of sampling directly in transformed spaces need further exploration. In many cases, it may be simpler to sample in the original space and transform the results.
@@ 5.3 Future Directions
Our work opens up several promising avenues for future research:
  • Expressive Transformations: While we focused on analytically tractable mappings, exploring more expressive transformations, such as invertible neural networks, could offer greater flexibility and adaptability to complex data geometries. This data-driven approach could lead to significant performance improvements, albeit at increased computational cost.
  • Theoretical Insights into Diffusion Processes: Deeper theoretical study is needed to understand the interplay between transformations and diffusion processes. Such insights could guide the selection of optimal transformations, challenging the prevailing assumption that the same space should be used for both training and sampling.
  • Inverse Problem Solving: The reverse-time Itô lemma presents a promising direction for solving inverse problems. By leveraging the ability to sample in transformed spaces, we can potentially develop more effective methods for inverse problem solving in various domains, building upon recent work in diffusion-based inverse problems [19, 6].
  • Extension to Other Score-Based Methods: Extending the score change of variables framework to other score-based methods, including variational inference [24] and energy-based models [30], presents a compelling direction for future research. This could unlock new capabilities within these frameworks and broaden their applicability.
  • Adaptive Transformation Selection: Developing methods to adaptively select or learn optimal transformations during the training or sampling process could significantly enhance the flexibility and performance of score-based models across diverse applications.
In summary, the theoretical foundations established here open the door to more versatile and effective methodologies in both training of score models and sampling from generative diffusion models. We believe the change of variables perspective will advance the performance, stability, and interpretability of score-based methods across a wide range of applications.